OpenStudy (anonymous):

The height , in feet, of an airborne tee-shirt t seconds after being launched can be approximated by h(t)=-15t^2+135t+10, 0

OpenStudy (anonymous):

-15t^2+135t+10=280 t^2-9t+18=0 (t-3)(t-6)=0 t=3or t=6 they are both can get ! why ? Do you write right?